Al's father is 45 years old.He is 15 years older than twice Al's age. how old is Al?

Found 2 solutions by FrankM, addingup:
Answer by FrankM(1040)

45 = 2A + 15
30 = 2A
A = 15

Answer by addingup(3677)

2A+15 = 45
2A = 30
A = 15
---------------------------------------
Check
2*15+15= 45
30+15= 45 We have the correct answer